OpenStudy (anonymous):

Find the coefficient x^-6 in the expansion of (3/2x)^12

OpenStudy (anonymous):

Can someone please help? Honestly have no idea what the hell I'm doing.

OpenStudy (anonymous):

i think there is a typo here

OpenStudy (anonymous):

nope, there is none

OpenStudy (anonymous):

i can do the equation for you if you want

OpenStudy (anonymous):

oh wait you were right

OpenStudy (anonymous):

imagine

OpenStudy (anonymous):

it's supposed to be "Find the coefficient of x^-6 in the expansion of \[(2x - 3/x^2)^{12}\]

OpenStudy (anonymous):

@satellite73 please don't leave me now man I really need your help

OpenStudy (anonymous):

ooh ok htat makes more sense

OpenStudy (anonymous):

ok so how do we go about solving this?

OpenStudy (anonymous):

\[(2x-\frac{3}{x^2})^{12}\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is going to be a bit of a pain, but we can do it what class?

OpenStudy (anonymous):

IB SL math

OpenStudy (anonymous):

and ya, I know, it's gonna be long and annoying ;-;

OpenStudy (anonymous):

anyway can we start?

OpenStudy (anonymous):

real long and real annoying first lets say what the answer is

OpenStudy (anonymous):

\[43,110,144\]

OpenStudy (anonymous):

ok now we can figure out how to do it

OpenStudy (anonymous):

the back of the book says the answer is \[\left(\begin{matrix}12 \\ 6\end{matrix}\right)\] \[* 2^6 * (-3)^6\]

OpenStudy (anonymous):

what terms will give \(x^{-6}\) that is a cheating answer

OpenStudy (anonymous):

but that is ok, maybe we can just figure out how they got that

OpenStudy (anonymous):

yeah, IB is really weird. anyway, idk what terms would find that

OpenStudy (anonymous):

when you expand , all the terms will look like \[\binom{12}{k}(2x)^k(-\frac{3}{x^2})^{n-k}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not sure if that makes any sense clear or no?

OpenStudy (anonymous):

Wait I don't think that's right? it should be (2x)^n-r

OpenStudy (anonymous):

am i wrong?

OpenStudy (anonymous):

lol don't be a slave to notation dear, your book uses \(r\) i guess, i used \(k\) same difference as the kids say

OpenStudy (anonymous):

no no no I mean you switched the order

OpenStudy (anonymous):

guess your book writes it as \[\binom{12}{r}(2x)^r(-\frac{3}{x^2})^{n-r}\]

OpenStudy (anonymous):

no, here I'll show you what they have

OpenStudy (anonymous):

oh i see what what you mean no need to write it

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

they must write \[\binom{12}{r}(2x)^{n-r}(-\frac{3}{x^2})^{r}\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I tried to do the shortcut way and plug in 6 but the negative is what's throwing me off :'(

OpenStudy (anonymous):

ok a couple typos there, nvm

OpenStudy (anonymous):

mkay

OpenStudy (anonymous):

so the real question is, what will give \(x^{-6}\) i.e. what is \(r\) in this case?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

or x/6?

OpenStudy (anonymous):

you have powers of \(x^2\) in the denominator and powers of \(x\) in the numerator

OpenStudy (anonymous):

sorry man I'm lost

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

each term will look something like \[C\frac{(2x)^k}{(-3x^2)^{12-k}}\]

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

like for example \[C\frac{(2x)^7}{(-3x^2)^{5}}\]

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

is that clear or no?

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

so the question is, when will the exponent just be \(-6\)

OpenStudy (anonymous):

in the example i did above, the exponent in the numerator is \(7\) and in the denominator it is \(10\) so all in all you will have \(x^{-3}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now that we see that (i hope) we can go ahead and figure out what \(r\) is

OpenStudy (anonymous):

hehe yea i got it

OpenStudy (anonymous):

ok so we have to find what would equal -6?

OpenStudy (anonymous):

no what i wrote was wrong sorry

OpenStudy (anonymous):

ok no biggie

OpenStudy (anonymous):

\[\frac{x^r}{x^{2({12-r)}}}=x^{-6}\]

OpenStudy (anonymous):

so..we still need to find the numbers or whatever that'll equal -6 right?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

soooo do we pick random numbers or is there something else we can do?

OpenStudy (anonymous):

i actually did not use algebra, i used trial and error, because it is pretty easy to see that \(r=-6\) because \[\frac{x^6}{x^{12}}=x^{-6}\]

OpenStudy (anonymous):

but we would use algebra if you needed to

OpenStudy (anonymous):

I mean I just wanna know what you plugged into r to make that happen. cos I know obviously that 6-12 = -6 but

OpenStudy (anonymous):

but I thought that wasn't the formula we were using

OpenStudy (anonymous):

yes that is what i plugged in and now that we know \(r=-6\) the answer they wrote should be clear you want to use algebra?

OpenStudy (anonymous):

just lemme look at it for a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

take your time, i will get a drink

OpenStudy (anonymous):

so you plugged in -6 for the r's?

OpenStudy (anonymous):

actually ya I wanna see the algebraic way. I know you did the easy way but I don't get how you're doing that

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

first, forgetting the coefficents, (because we are just trying to find \(r\)) we know the terms look like \[\frac{x^r}{(x^2)^{12-r}}\] or \[\huge \frac{x^r}{x^{24-2r}}\] right?

OpenStudy (anonymous):

yessir

OpenStudy (anonymous):

if we were a bit more sophisticated we would have subtracted the exponents already (i was only trying to explain so it made sense) and say that each term looks like \[\huge x^{3r-24}\]

OpenStudy (anonymous):

yeah I get that

OpenStudy (anonymous):

hope you see that i subtracted the exponents as always when you divide, and \[r-(24-2r)=3r-24\]

OpenStudy (anonymous):

ya lol

OpenStudy (anonymous):

you want that to equal \(-6\) so set \[3r-24=-6\] and solve for \(r\)

OpenStudy (anonymous):

you get \(r=6\) pretty much in your head

OpenStudy (anonymous):

6

OpenStudy (anonymous):

omg wait so we just found the general term

OpenStudy (anonymous):

i mean we found r but we needed to find the general term first to do that

OpenStudy (anonymous):

right now that is NOT how i did it, i just wrote it down on paper, and saw that 6 worked to give \(-6\) as the exponents, but that is the algebra that does it

OpenStudy (anonymous):

ok so that helped a ton. thanks. but how did they get the answer they got?

OpenStudy (anonymous):

i mean i understand everything of the answer up until the (-3)^6 part

OpenStudy (anonymous):

ok now we know \(r=6\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we know the binomial coefficient will be \[\binom{12}{6}\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and we don't have \(x\) in the numerator, we have \(2x\) so if \(r=6\) and of course that makes \(12-r=6\) also, you have \[\binom{12}{6}2^6(-3)^6x^{-6}\]

OpenStudy (anonymous):

umm..gimme a second

OpenStudy (anonymous):

just to look at it

OpenStudy (anonymous):

if it was \[(x+\frac{1}{x^2})^{12}\]THEN we would have \[\binom{12}{6}x^{-6}\]